Required information Problem 11-4A Analysis of changes in stockholders' equity accounts LO C3, P2, P3 [The following information applies to the questions displayed below.] The equity sections from Atticus Group’s 2016 and 2017 year-end balance sheets follow. Stockholders’ Equity (December 31, 2016) Common stock—$6 par value, 100,000 shares authorized, 30,000 shares issued and outstanding $ 180,000 Paid-in capital in excess of par value, common stock 140,000 Retained earnings 360,000 Total stockholders’ equity $ 680,000 Stockholders’ Equity (December 31, 2017) Common stock—$6 par value, 100,000 shares authorized, 35,200 shares issued, 4,000 shares in treasury $ 211,200 Paid-in capital in excess of par value, common stock 171,200 Retained earnings ($60,000 restricted by treasury stock) 420,000 802,400 Less cost of treasury stock (60,000 ) Total stockholders’ equity $ 742,400 The following transactions and events affected its equity during year 2017. Jan. 5 Declared a $0.50 per share cash dividend, date of record January 10. Mar. 20 Purchased treasury stock for cash. Apr. 5 Declared a $0.50 per share cash dividend, date of record April 10. July 5 Declared a $0.50 per share cash dividend, date of record July 10. July 31 Declared a 20% stock dividend when the stock’s market value was $12 per share. Aug. 14 Issued the stock dividend that was declared on July 31. Oct. 5 Declared a $0.50 per share cash dividend, date of record October 10. Problem 11-4A Part 5 5. How much net income did the company earn during year 2017?
Answer #1

Calculation of Cash Dividend

Date Common Stock Dividend
Jan 5 (30000*$0.50)= $15000
Apr 5 (30000-4000)*$0.50= 13000
July 5 (26000*$0.50)= 13000
Oct 5 (26000*1.20*$0.50)= 15600
Total cash dividend $56600

Calculation of Stock Dividend

Stock dividend= (30000-4000)*20%*$12= $62400

Calculation of Net Income

Ending Retained earnings= Beginning Retained earnings+Net income-Cash dividend-Stock dividend

420000= $360000+Net income-56600-62400

$420000-360000+56600+62400= Net income

Net income= $179000